OpenStudy (anonymous):

What is the equation for an ellipse with foci at (0, 8) and (0, -8) with co-vertices (8, 0) and (-8, 0) ?

OpenStudy (mertsj):

If the vertices are (8,0) and (-8,0) then the center is the origin and the major axis is horizontal. Would you agree?

OpenStudy (anonymous):

no because the vertices would be vertical, the co-vertices are horizontal

OpenStudy (mertsj):

Oh. my bad. The covertices are (8,0) and (-8,0). So it is vertical and its equation will be of the form: \[\frac{x^2}{b^2}+\frac{y^2}{a^2}=1\]

OpenStudy (mertsj):

So all we have to do is recognize that b = 8 and c= 8. Use a_2-b^2=c^2 and find a.

OpenStudy (anonymous):

 This is what I keep getting but I don't think its right because the properties would be different. \[x ^{2}+ \frac{ y ^{2} }{ 64 } =1\]